OpenStudy (anonymous):

Solve using the substitution method. Show work please. 7. x= -2y x= 2-4y 8. r= -3s r= 10 -4s 9. x= 3y-4 2x-y=7 10. s+t =-4 s-t=2 11. x-y=6 x+y= -2 12. y-2x= -6 2y-x =5

OpenStudy (anonymous):

(7) x=-2y, x=2-4y right?

OpenStudy (anonymous):

right

OpenStudy (texaschic101):

7. x = -2y x = 2 - 4y we know that x = -2y, so sub -2y in for x in the 2nd equation... x = 2 - 4y (-2y) = 2 - 4y (add 4y to each side) -2y + 4y = 2 - 4y + 4y -2y = 2 (divide by -2 to get x by itself) -2y/-2 = 2/-2 y = -1 now sub -1 in for y in the first equation... x = -2y x = -2(-1) x = 2 check... x = -2y 2 = -2(-1) 2 = 2 (correct) I just did one to show you how to do it..... Get one of the variables and substitute it into the other equation to get the answer to one variable. Then substitute the variable known into the other equation to find the unknown variable. You can check by substituting both known variables into either of the original equations :)

OpenStudy (anonymous):

Thank you very much I now understand :)

OpenStudy (texaschic101):

That is good....that is what I was going for :)